OpenStudy (anonymous):

If the sin 60° is square root 3 over 2, then the cos ____° = _____. 30° ; 1 over 2 60° ; square root 2 over 2 30° ; square root 3 over 2 60° ; 1

OpenStudy (anonymous):

I think it is either A or C because this is for a right triangle and if one angle is 60 the other must be 90 because of the right angle. which concludes the last angle being 30 but I don't understand the fraction part.

OpenStudy (whpalmer4):

Sin and Cos are complementary in that \[\sin x^\circ = \cos (90^\circ - x^\circ)\]

OpenStudy (anonymous):

im confused

OpenStudy (whpalmer4):

okay. what is confusing you?

OpenStudy (anonymous):

how does sin = cos sin is opposite over hypetunuse and cos is adjacent over hyputenuse

OpenStudy (whpalmer4):

I didn't say that sin = cos, I said that sin x = cos (90-x) and in fact, there are values of x where sin x = cos x. Draw yourself a 45/45/90 triangle and you'll see one.

OpenStudy (anonymous):

o

OpenStudy (whpalmer4):

do you have a diagram of the unit circle?